OpenStudy (anonymous):

( 7 + 2 ) ex2 - 2ex3=

OpenStudy (anonymous):

medal @babyboy2332 @cmtboy2

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

I don't get what ex2 ,ex3 mean

OpenStudy (anonymous):

exponent @Loser66

OpenStudy (michele_laino):

I think that they are two function like this: \[\huge {e^{{x^2}}},\quad {e^{{x^3}}}\]

OpenStudy (michele_laino):

functions*

OpenStudy (michele_laino):

am I right?

OpenStudy (loser66):

@Michele_Laino it is Middle high, not college stuff. I think it is jus (7+2)^2 -2^3

OpenStudy (loser66):

Am I right? @alibaby

OpenStudy (anonymous):

yes

OpenStudy (loser66):

ok, 7+2 =?

OpenStudy (anonymous):

9 @Loser66

zepdrix (zepdrix):

\[\large\rm (7+2)^2 -2^3\]\[\large\rm =9^2 -2^3\]Good :) that takes care of the first step. Understand how to deal with the exponents now?

OpenStudy (anonymous):

yes